OpenStudy (anonymous):

Given: 2x+3y=12 2x-y=4 What is the value of x + y? Please explain! Thanxs!!!

OpenStudy (anonymous):

Weird question at first sight but here goes. Subtract equation 2 from equation 1 giving 4y = 8 Here's why, 2x - 2x = 0, 3y - (-y) = 4y and 12 -4 = 8 hence 4y = 8 From this we get y = 2 by dividing both sides by 4. Substitute the value of y ( 2 remember) into second equation to discover value x. 2x - 2 = 4 which implies (by adding 2 to both sides) 2x = 6 and hence x = 3 Therefore the answer to x+ y is 2 +3 = 5